OpenStudy (anonymous):

f(x) square root x+5 and g(x) x^2-x what is (f+g)(11) ??

OpenStudy (anonymous):

(f+g)11=f(11)+g(11)

myininaya (myininaya):

\[(f+g)(x)=\sqrt{x+5}+(x^2-x)\] so just replace x with 11 and evaluate and simplify